Which of the following quadratic equations has solutions x=6a and x=-3b?
A. x^2-18ab=0
B. x^2-x(3b-6a)-18ab=0
C. x^2-x(3b+6a)+18ab=0
D. x^2+x(3b-6a)-18ab=0
E. x^2+x(3b+6a)+18ab=0

Answer by stanbon(75887)

Which of the following quadratic equations has solutions x=6a and x=-3b?
---
Modify:
x-6a = 0
x+3b = 0
---
Multiply:
x^2 -6ax+3bx-18ab = 0
Rearrange:
x^2 -x(6a-3b)-18ab = 0
x^2+x(3b-6a)-18ab = 0
---
Ans D
==========
Cheers,
Stan H.
==========
A. x^2-18ab=0
B. x^2-x(3b-6a)-18ab=0
C. x^2-x(3b+6a)+18ab=0
D. x^2+x(3b-6a)-18ab=0
E. x^2+x(3b+6a)+18ab=0